OpenStudy (anonymous):

Simplify the given expression. sqaure root -108 + sqaure root -27. Please help, cause I'm so lost..

OpenStudy (earthcitizen):

\[\sqrt{-108}+\sqrt{-27}= (\sqrt{108}*i)+(\sqrt{27}*i)\]

OpenStudy (anonymous):

it should be 6isquareroot6+3isquareroot3

OpenStudy (anonymous):

you cant add radicals that arent the same

OpenStudy (earthcitizen):

\[9i \sqrt{3}\]

OpenStudy (anonymous):

earth citizen is correct i looked at my paper wrong sorry

OpenStudy (turingtest):

do you understand the answer?

OpenStudy (anonymous):

Me? Not really.

OpenStudy (turingtest):

Do you want to?

OpenStudy (anonymous):

Yes

OpenStudy (turingtest):

ok, look at just the first term:\[\sqrt{-108}\]factor the number under the radical sign:\[\sqrt{-1\times2\times2\times3\times3\times3}\]we know that \[\sqrt{-1}=i\]so we can pull that out of the radical. Also, whenever we see two factors the same under the radical that means we can pull out one copy:\[\sqrt{-1\times2\times2\times3\times3\times3}=\sqrt{2^2\times3^2\times(-1)\times3}=(2\times3)i\sqrt3\]\[=6i\sqrt3\]...

OpenStudy (turingtest):

you can do the same with \[\sqrt{-27}=\sqrt{-1\times3\times3\times3}=\sqrt{-1\times3^2\times3}=3i\sqrt3\]adding the two together then gives\[\sqrt{-108}+\sqrt{-27}=6i\sqrt3+3i\sqrt3=9i\sqrt3\]I hope that help you understand the problem better.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (turingtest):

welcome :)

OpenStudy (earthcitizen):

thnx!